OpenStudy (anonymous):

If you rent a car for one day and drive it for 100 miles, the cost is $40.00. If you drive it 220 miles the cost is $46.

OpenStudy (ghazi):

so what is asked in this question?

OpenStudy (anonymous):

Let c be the cost per day to rent the car and m be the cost per mile that the car is driven. Solve the following for c and m:\[\{c+100m=40, c+220m=46\} \]\[\left\{c=35,m=\frac{1}{20}\right\} \text{ or } \{c=\text{$\$$35} \text{ per} \text{ day},m= \text{one} \text{ nickle} \text{ per} \text{ mile}\} \]